OpenStudy (jennychan12):

Evaluate the limit (see below) over the region bounded by the graphs of the equations. (see below)

OpenStudy (jennychan12):

\[\lim_{n \rightarrow \infty} \sum_{i =1}^{n} f(c _{1}) \Delta x _{i}\] for \[f(x) = \sqrt{x}\] y = 0 x = 0 x = 2

OpenStudy (abb0t):

So, this is a Reimmann Sum Qestion essentially. And you know that \[\sqrt{x} = x^\frac{ 1 }{ 2 }\]

OpenStudy (jennychan12):

mhm

OpenStudy (abb0t):

I didn't use antiderivs :P \[\int\limits_{a}^{b}f(x)dx = \lim_{n \rightarrow ∞}\sum_{i=1}^{n}f(x_i^*)\Delta x\]Reimann sum's is basiclaly the same thing as:

OpenStudy (zehanz):

If c_1 is supposed to mean x_i, then you have:\[\lim_{n \rightarrow \infty}\sum_{i=1}^{n} \sqrt{x_i}\Delta x_i:=\int\limits_{0}^{2}\sqrt{x}dx\]It is the definition of the integral.

OpenStudy (abb0t):

\[\Delta x = \frac{ b-a }{ n }\]

OpenStudy (abb0t):

n is a constant while i changes.

OpenStudy (jennychan12):

i know it's the integral, but i think they want us to use the riemann sum with the xi stuff. if it is like that, then i just choose a random n? the attachment is an example.

OpenStudy (abb0t):

You don't choose a random n, you simply keep n there. For instance, if you have 0<x<2, then ∆x = 2-0/n = 2/n

OpenStudy (jennychan12):

ohh that...

OpenStudy (jennychan12):

but what's the y = 0 for? is it just saying it's above the x axis?

OpenStudy (zehanz):

They surely make your life difficult! The whole reasoning behind Riemann sums is: at the moment it is not meant as an approximation anymore, but the real thing (i.e. lim n to infty), the Fundamental Theorem of Calculus comes to help and we "just" have to calc the difference of two values of a primitive function.

OpenStudy (abb0t):

Well of course, the graph of √x starts at 0 and goes up. You can't have a neg number under the radical. then u'd be dealing w/ imaginary numbers, which isn't pleasant.

OpenStudy (jennychan12):

oh haha ok.

OpenStudy (abb0t):

Useful formulas to know for this problem: \[\sum_{i=1}^{n}c = cn\] \[\sum_{i=1}^{n}i = \frac{ n(n+1) }{ 2 }\] \[\sum_{i=1}^{n}i^2 = \frac{ n(n+1)(2n+1) }{ 6 }\] \[\sum_{i=1}^{n}i^3 = \left[ \frac{ n(n+1) }{ 2 } \right]^2\]

OpenStudy (jennychan12):

so then the xi = 2/n? and then u plug it into √x to get √(2i/n) and you just take the limit as n--> ∞ ? oh i forgot. it also says (HInt: Let \[ c_{i} = \frac{ 2i^2 }{ n }\] with that, does that mean i just use that as xi?

OpenStudy (abb0t):

As you cycle through the integers from 1 to n in the summation only "i" changes and so anything that isn’t an "i" will be a constant and can be factored out of the summation. So you will be using: the 3rd formula I provided and taking the limit of that as n->∞

OpenStudy (jennychan12):

yeah i know the formulas, thanks. okay thanks :)

OpenStudy (abb0t):

Welcome. If you get stuck, let me know :) But it sounds to me like you understand it. Good luck.

OpenStudy (jennychan12):

thanks. yeah i kinda forgot about this after learning integrals. :/